1. A beam of light is incident, from air, on a flat surface of glass (n=1.56). The refracted ray inside the glass is found to make an angel of 27.7 ° with the direction of the incident ray. Calculate the angle of incident.
Answer #1

Here = 1.56 na cair) & r = 27.7° So by Shells law. nisini =n2 sina 1x ini: 1.56 X Gn2-7-7 Sini=0-7251S i=46-48 o I=46-5